Consider the overall reaction below, which is experimentally observed to be second order in AB and zero order in C.

AB+ C ---------> A+BC

Is the following mechanism valid for this reaction? State the two key requirements and weather they are followed in this case .

AB + AB -----> AB2 + A (slow)

AB2 + C ------> AB + BC (fast)
Answer #1

Solution :

The mechanism of the reaction is valid.

The reaction is as follows:

AB + C  = A + BC

The reaction is second order with AB and zero order with respect to C, therefore, the experimental rate law expression for the reaction is as follows:

Rate = k [AB]^2 [C]

The reaction mechanism is as follows:

AB + AB = AB2 + A (slow step)

AB2 + C = AB + BC (fast step)

After adding two steps,

2AB + C = AB + BC

The rate depends on the slowest step. Therefore, the observed rate law is as follows:

Rate = k [AB]^2

The observed rate law is same as the experimental rate law. Hence, mechanism is valid.

The two key requirements are:

1)The slowest elementary step matches the given rate law.

2)The mechanisms matches the given reaction after crossing out the intermediates of the reactions in the mechanisms.